# Membership Inference Attacks on Recommender System: A Survey

### v1

@@ I Introduction
Recommendation systems (RecSys) have seen significant advances over the past decade and have been widely used in various scenarios such as job matching [6], e-commerce [5], and entertainment [8] and social media [15]. Besides the advanced algorithm design and powerful computational resources, the availability of large datasets is another key factor contributing to the success of RecSys [31]. As datasets can contain individuals’ private information such as user purchase records, connected friends, and watched movies, it is essential that RecSys model owners should not leak privacy sensitive information about their training data. However, recent studies [29, 28, 33, 25, 7, 14, 12, 32]have shown that RecSys models are prone to memorizing information of user data, making them vulnerable to membership inference attacks [17, 4, 3, 9, 11, 22, 10]. Membership inference attacks (MIAs), which are also the focus of this article, aim to infer members of the training dataset of an ML model. We discuss MIAs in detail subsequently.
@@
MIAs on RecSys models aim to infer whether a user record was used to train a RecSys model or not. MIAs can raise severe privacy risks to individuals. For example, by identifying the fact that purchase records have been used to train the E-commerce RecSys model associated with a certain user, MIAs can infer that user’s special quirks. A recent report [23] published by the National Institute of Standards and Technology (NIST) specifically mentions that an MIA determining an individual was included in the dataset used to train the target model is a confidentiality violation. Moreover, such privacy risks caused by MIAs can lead to commercial companies who provide product services through RecSys to their users violating privacy regulations such as General Data Protection Regulation (GDPR) [24], CCPA [1].
@@
Membership Inference Attacks (MIAs) on machine learning (ML) were first proposed by Shokri et al. [22], who demonstrated that an attacker can determine whether a data record was used to train an ML model solely from the prediction vector of that record (i.e., with black-box access to the target model). Since then, a growing number of studies have investigated MIAs across various domains, including computer vision [3], natural language processing [20, 18, 26], and audio processing [21]. Research on MIAs in RecSys has emerged relatively later than in other domains, yet it remains highly important. User privacy and security directly shape users’ trust in the system, which in turn influences a company’s reputation. The first work to introduce MIAs in RecSys was conducted by Zhang et al. [29]. However, traditional MIAs are difficult to apply to RecSys because they typically rely on posterior probabilities to construct the attack model. In contrast, adversaries in RecSys cannot observe posterior probabilities; they only have access to the ranked item lists output by the model. Furthermore, the ranking property of these lists introduces an additional layer of complexity. Designing MIAs that can effectively leverage and adapt to these characteristics remains one of the key challenges in this research area. Subsequently, a large body of work has proposed different MIAs for RecSys, which can generally be divided into two categories: user-level and user–interaction-level attacks. Given the importance of user privacy protection and the widespread deployment of RecSys, both academia and industry have shown strong interest in this research area. In this article, we present the first systematic study that summarizes different membership inference attacks and defenses on RecSys and establishes taxonomies based on multiple criteria for the research community. Unlike existing reviews that include only a very limited number of publications on MIAs in RecSys—for example, the survey in [29] cites only a single relevant work—we conduct a comprehensive search and include nearly all related studies in this survey. Our survey provides a deeper discussion of the underlying concepts, theories, methods, and taxonomies, as well as visions for future research directions. The main contributions of this article are as follows:
@@
The rest of the article is organized as follows: Section II introduces MIAs on RecSys preliminaries. Section III introduces the existing attack approaches and provide taxonomies to categorize the released papers. In Section 4, we discuss current defense on MIAs-RecSys. discuss the challenges and propose the future directions. concludes this article.
II Preliminary
@@
Recommender System (RecSys) Recommender systems (RecSys) have become a fundamental component of many online services, ranging from e-commerce and social media platforms to digital entertainment and news portals. By analyzing user–item interactions[16, 15], RecSys are able to predict user preferences and provide personalized recommendations, thereby improving user experience and engagement. With the increasing scale and ubiquity of RecSys in real-world applications, ensuring their reliability, fairness, and privacy has become an important research focus. In particular, user data in RecSys often contain sensitive information, making these systems potential targets for privacy attacks, such as membership inference attacks (MIAs). In the following section, we discuss MIAs in RecSys depthly.
III Memebership Inference Attacks on RecSys
III-A user-level MIAs
@@ III Memebership Inference Attacks on RecSys
Ming et al.[29] were the first to propose a user-level membership inference attack in RecSys based on item embedding differences, aiming to infer whether a user’s data was included in the training set of a recommender model. Under the black-box setting, the attacker can only observe the output recommendation list. The item-embedding-based method assumes that if a user’s data is included in the training set, the recommended items should be similar to the items the user has interacted with. In contrast, for non-member users, the system lacks knowledge of their preferences and therefore cannot generate personalized recommendations. In the original work, the top popular items were used as the recommendations for non-members. To construct features for attack, the embeddings of the interacted items were averaged, and the embeddings of the recommended items were averaged; the difference between these two vectors was then used as the membership feature. A two-layer multilayer perceptron (MLP) was trained as the attack model to distinguish members from non-members. However, this approach is unstable when the target dataset and the shadow dataset differ. The reason is that the method implicitly assumes that item embeddings are generated by a fixed collaborative filtering model. In practice, embeddings trained by different algorithms (e.g., BPR vs. NeuMF) can diverge in the latent space even when trained on the same dataset. Furthermore, the item embeddings generated by the target and shadow models may also exhibit discrepancies, which degrades attack performance. To address this issue, Wang et al. [25] proposed a debiased learning MIA (DL-MIA) framework for RecSys that mitigates the embedding bias between the target and shadow models. For DL-MIA, to overcome the limitations of Item Difference MIAs (ID-MIA), which suffer from (i) a training-data bias—distributional gaps between shadow and target recommenders that make attack samples generated by the shadow model poorly transferable—and (ii) an estimation bias—since the attacker cannot observe hidden states (user/item embeddings), externally constructed difference vectors are noisy—the DL-MIA framework explicitly debiases learning within the difference-vector paradigm. Concretely, it (1) builds difference vectors from each user’s history and the system’s recommendation list; (2) employs a variational auto-encoder–based disentangled encoder to separate recommender-invariant from recommender-specific features, narrowing the shadow–target gap and mitigating training-data bias; (3) learns a truth-level score per difference vector as a sample weight to discount poorly estimated features, thereby mitigating estimation bias; and (4) trains a member/non-member classifier (MLP) on the disentangled, reweighted representations, optimized with an alternating training procedure. Empirically, DL-MIA simultaneously reduces both biases and achieves state-of-the-art attack performance across general and sequential recommenders.
@@
Although subsequent experiments demonstrated that DL-MIA is more accurate and stable than the ID-MIA under target–shadow mismatches, its effectiveness remains sensitive to the information exposed to the adversary—most notably the dimensionality of item embeddings and the length of the top– recommendation list. Empirically, performance tends to improve with larger embedding dimensions and larger , whereas short lists (e.g., ) often provide a weak signal in practice. Moreover, obtaining an in–distribution shadow dataset is unrealistic, and the common assumption that non–member recommendations align with a globally popular item set is brittle. To address both the data–distribution limitation and the member/non–member mode gap, Zhu et al.[33] propose Model Extraction based MIA (ME-MIA) for sequential recommenders, which first extracts a surrogate model via black–box queries that aligns its recommended item set and rank order with the target using generic list–consistency objectives, and then performs membership inference using the surrogate’s richer signals (scores, ranks, similarities).[29, 25, 33] ME-MIA targets sequential recommenders in a black-box setting by first extracting a surrogate model that imitates the target’s top – items and their rank order using two generic objectives—ranking consistency and positive-item consistency—derived solely from recommendation lists. Because the surrogate’s training does not reveal membership labels, a shadow model (trained with the same objective) is used to construct labeled data; rich signals from the surrogate (scores, ranks, similarities) then feed a binary classifier for membership inference. To reduce data demands, ME-MIA offers (i) a data-efficient variant that augments sequences by replacing actions with nearest-neighbor items in the surrogate’s embedding space, and (ii) a data-free variant that synthesizes sequences autoregressively and queries the target for soft labels, following data-free model-extraction practice. These steps yield effective and transferable attacks under black-box sequential settings. While ME-MIA demonstrates that membership inference is possible even in data–free settings, this comes with a loss in effectiveness; moreover, the role of item scoring in shaping attack success remains underexplored, and it is unclear whether ME-MIA readily generalizes to non-sequential RecSys. Despite recent progress at the user level, open challenges persist: how to design more efficient attacks, reduce confounding factors that obscure the signal (e.g., embedding dimensionality, top – , popularity effects), and accommodate the inherent heterogeneity of recommender systems with model–agnostic methods. Beyond traditional RecSys MIAs, He et al.[14] (to our knowledge) present the first user–level attacks for LLM–based recommender systems, proposing five attacks—Inquiry, Hallucination, Semantic, Poisoning, and Confusion—that exploit LLMs’ memorization, text generation, and hallucination behaviors; evaluated across multiple popular LLMs, these attacks demonstrate clear effectiveness in revealing whether specific user were included in in–context prompts.
III-B Interaction-level MIAs
@@
RecSys pose privacy risks beyond the user level: in addition to whether a user appears in the training set, one may care about the presence of a specific user–item interaction (interaction-level privacy). Yuan et al. [28] conducted the first systematic study of interaction-level membership inference in federated recommender systems (FedRec), showing that such attacks are feasible and effective. In contrast to traditional FL-MIAs which have largely been designed for classification and aim to infer whether a sample was used for training (and sometimes which client used it), FedRecs are built on collaborative filtering, where item embedding dynamics can leak whether a client’s item set has changed. Moreover, each FedRec client corresponds to an individual user whose local data are non-IID, and the local dataset naturally consists of interacted and non-interacted items (only the former are meaningful targets for interaction-level MIAs), making the shadow-data assumptions commonly used in FL-MIAs impractical. Finally, unlike many FL classification settings that conceptually “share” a single global parameter set, FedRecs typically keep user embeddings private on the client while sharing (or aggregating) item embeddings on the server—an architectural choice that changes the attack surface and defenses. For these challenges, Yuan et al [28] proposed a white-box interaction-level MIA on FedRec (c). IFed-MIA supposes the adversary know the target user’s updated public parameter (including gradient) and some hyperparameter like learning rate and negative sampling ratio. IFed-MIA observes only clients’ uploaded item-embedding updates and public hyperparameters (e.g., learning rate and negative-sampling ratio), but not private user embeddings or raw histories. From these uploads, the server can tell which items participated in a user’s local training; the problem then reduces to deciding whether a target user–item pair is a positive interaction. The attack relies on an empirical distance test: training three local models— on , on with a different private initialization, and on where is flipped—yields , enabling accurate inference of . Operationally, the server (i) identifies updated items, (ii) constructs candidate labels for them, (iii) trains a fake local model on the constructed dataset, and (iv) selects positives by comparing item-embedding distances between the uploaded parameters and the fake model, iterating until a prescribed ratio is met [28]. In practice, training hyperparameters such as the negative sampling ratio are typically not observable to an external adversary, casting doubt on the practicality of attacks that rely on them; even in federated recommenders, such hyperparameters are assumed known only to an honest-but-curious server that broadcasts them to clients. To improve interaction-level attacks beyond these assumptions, Zhong et al.[32] propose a knowledge-graph (KG)–based method (MINER) that constructs a similarity matrix between recommended and interacted items using multiple distance measures and employs a cumulative-learning bilateral-branch network to emphasize long-tail items, thereby heightening risks tied to personalized interactions. However, while MINER targets the long-tail, it shows limited effectiveness in the low-FPR regime and its reliance on KGs constrains applicability to other RecSys architectures. Building on these observations, He et al.[12] design a LiRA-based interaction-level attack (RecLiRA) that attains state-of-the-art true-positive rates at low false-positive rates across common RecSys models; moreover, they motivate a differential-privacy–inspired privacy score, , to quantify interaction-level risk and aggregate it to the user level[13]. Also, RecLiRA can be applied to all recommendation systems with implicit feedback. He et al.[14] pioneer interaction–level MIAs for LLM–based recommender systems, exploiting LLMs’ tendency to memorize prompt content through direct inquiry and hallucination (“confusion”) attacks; experiments on several widely used LLMs and benchmark datasets show strong attack advantages, underscoring practical risks for in–context LLM RecSys.
III-C Evaluation Metric

### v2

@@ 1. Introduction
Recommendation systems (RecSys) have seen significant advances over the past decade and have been widely used in various scenarios such as job matching (Chen et al., 2023a), e-commerce (Chen et al., 2023b), and entertainment (Gao et al., 2023) and social media (He et al., 2020). Besides the advanced algorithm design and powerful computational resources, the availability of large datasets is another key factor contributing to the success of RecSys (Zhao et al., 2024). As datasets can contain individuals’ private information such as user purchase records, connected friends, and watched movies, it is essential that RecSys model owners should not leak privacy sensitive information about their training data. However, recent studies (Zhang et al., 2021; Yuan et al., 2023; Zhu et al., 2023; Wang et al., 2022; Chi et al., 2024; He et al., 2025c, a; Zhong et al., 2024)have shown that RecSys models are prone to memorizing information of user data, making them vulnerable to membership inference attacks (Hu et al., 2022; Carlini et al., 2022b, a; Gu et al., 2024; Hayes et al., [n. d.]; Shokri et al., 2017; Gu et al., 2025). Membership inference attacks (MIAs), which are also the focus of this article, aim to infer members of the training dataset of an ML model. We discuss MIAs in detail subsequently.
@@
MIAs on RecSys models aim to infer whether a user record was used to train a RecSys model or not. MIAs can raise severe privacy risks to individuals. For example, by identifying the fact that purchase records have been used to train the E-commerce RecSys model associated with a certain user, MIAs can infer that user’s special quirks. A recent report (Tabassi et al., 2019) published by the National Institute of Standards and Technology (NIST) specifically mentions that an MIA determining an individual was included in the dataset used to train the target model is a confidentiality violation. Moreover, such privacy risks caused by MIAs can lead to commercial companies who provide product services through RecSys to their users violating privacy regulations such as General Data Protection Regulation (GDPR) (Voigt and Von dem Bussche, 2017), CCPA (ccp, 2018).
@@
Membership Inference Attacks (MIAs) on machine learning (ML) were first proposed by Shokri et al. (Shokri et al., 2017), who demonstrated that an attacker can determine whether a data record was used to train an ML model solely from the prediction vector of that record (i.e., with black-box access to the target model). Since then, a growing number of studies have investigated MIAs across various domains, including computer vision (Carlini et al., 2022a), natural language processing (Nasr et al., 2019; Liu et al., 2022; Wen et al., 2024), and audio processing (Shah et al., 2021). Research on MIAs in RecSys has emerged relatively later than in other domains, yet it remains highly important. User privacy and security directly shape users’ trust in the system, which in turn influences a company’s reputation. The first work to introduce MIAs in RecSys was conducted by Zhang et al. (Zhang et al., 2021). However, traditional MIAs are difficult to apply to RecSys because they typically rely on posterior probabilities to construct the attack model. In contrast, adversaries in RecSys cannot observe posterior probabilities; they only have access to the ranked item lists output by the model. Furthermore, the ranking property of these lists introduces an additional layer of complexity. Designing MIAs that can effectively leverage and adapt to these characteristics remains one of the key challenges in this research area. Subsequently, a large body of work has proposed different MIAs for RecSys, which can generally be divided into two categories: user-level and user–interaction-level attacks. Given the importance of user privacy protection and the widespread deployment of RecSys, both academia and industry have shown strong interest in this research area. In this article, we present the first systematic study that summarizes different membership inference attacks and defenses on RecSys and establishes taxonomies based on multiple criteria for the research community. Unlike existing reviews that include only a very limited number of publications on MIAs in RecSys—for example, the survey in (Zhang et al., 2021) cites only a single relevant work—we conduct a comprehensive search and include nearly all related studies in this survey. Our survey provides a deeper discussion of the underlying concepts, theories, methods, and taxonomies, as well as visions for future research directions. The main contributions of this article are as follows:
@@
The rest of the article is organized as follows: Section 2 introduces MIAs on RecSys preliminaries. Section 3 introduces the existing attack approaches and provide taxonomies to categorize the released papers. In Section 4, we discuss current defense on MIAs-RecSys. discuss the challenges and propose the future directions. concludes this article.
2. Preliminary
@@
Recommender System (RecSys) Recommender systems (RecSys) have become a fundamental component of many online services, ranging from e-commerce and social media platforms to digital entertainment and news portals. By analyzing user–item interactions(He et al., 2017, 2020), RecSys are able to predict user preferences and provide personalized recommendations, thereby improving user experience and engagement. With the increasing scale and ubiquity of RecSys in real-world applications, ensuring their reliability, fairness, and privacy has become an important research focus. In particular, user data in RecSys often contain sensitive information, making these systems potential targets for privacy attacks, such as membership inference attacks (MIAs). In the following section, we discuss MIAs in RecSys depthly.
3. Memebership Inference Attacks on RecSys
3.1. user-level MIAs
@@ 3. Memebership Inference Attacks on RecSys
Ming et al.(Zhang et al., 2021) were the first to propose a user-level membership inference attack in RecSys based on item embedding differences, aiming to infer whether a user’s data was included in the training set of a recommender model. Under the black-box setting, the attacker can only observe the output recommendation list. The item-embedding-based method assumes that if a user’s data is included in the training set, the recommended items should be similar to the items the user has interacted with. In contrast, for non-member users, the system lacks knowledge of their preferences and therefore cannot generate personalized recommendations. In the original work, the top popular items were used as the recommendations for non-members. To construct features for attack, the embeddings of the interacted items were averaged, and the embeddings of the recommended items were averaged; the difference between these two vectors was then used as the membership feature. A two-layer multilayer perceptron (MLP) was trained as the attack model to distinguish members from non-members. However, this approach is unstable when the target dataset and the shadow dataset differ. The reason is that the method implicitly assumes that item embeddings are generated by a fixed collaborative filtering model. In practice, embeddings trained by different algorithms (e.g., BPR vs. NeuMF) can diverge in the latent space even when trained on the same dataset. Furthermore, the item embeddings generated by the target and shadow models may also exhibit discrepancies, which degrades attack performance. To address this issue, Wang et al. (Wang et al., 2022) proposed a debiased learning MIA (DL-MIA) framework for RecSys that mitigates the embedding bias between the target and shadow models. For DL-MIA, to overcome the limitations of Item Difference MIAs (ID-MIA), which suffer from (i) a training-data bias—distributional gaps between shadow and target recommenders that make attack samples generated by the shadow model poorly transferable—and (ii) an estimation bias—since the attacker cannot observe hidden states (user/item embeddings), externally constructed difference vectors are noisy—the DL-MIA framework explicitly debiases learning within the difference-vector paradigm. Concretely, it (1) builds difference vectors from each user’s history and the system’s recommendation list; (2) employs a variational auto-encoder–based disentangled encoder to separate recommender-invariant from recommender-specific features, narrowing the shadow–target gap and mitigating training-data bias; (3) learns a truth-level score per difference vector as a sample weight to discount poorly estimated features, thereby mitigating estimation bias; and (4) trains a member/non-member classifier (MLP) on the disentangled, reweighted representations, optimized with an alternating training procedure. Empirically, DL-MIA simultaneously reduces both biases and achieves state-of-the-art attack performance across general and sequential recommenders.
@@
Although subsequent experiments demonstrated that DL-MIA is more accurate and stable than the ID-MIA under target–shadow mismatches, its effectiveness remains sensitive to the information exposed to the adversary—most notably the dimensionality of item embeddings and the length of the top – recommendation list. Empirically, performance tends to improve with larger embedding dimensions and larger , whereas short lists (e.g., ) often provide a weak signal in practice. Moreover, obtaining the training dataset distribution is unrealistic. Chi et al(Chi et al., 2024). proposed the shadow-free MIA method to infer the membership without training the shadow model, compared with the previous methods(Zhang et al., 2021; Wang et al., 2022; Zhu et al., 2023). The intuition of the SF-MIA is to compare the recommended items with general popular items (which can be obtained by creating an empty account). If the recommended items align more closely to popular items, then it’s likely not a member. But if the recommended items have a higher similarity to the user’s historical interactions, then it’ll be classified as a member. The experiment results show that this shadow-free approach not only significantly reduces computational cost for MIAs, but also reaches a comparable performance to previous shadow-training approaches. At the same time, since there is no need to train the shadow model, this method reduces the time cost of the attack. However, the common assumption that non–member recommendations align with a globally popular item set is brittle in the real world. The current sequential RecSys can reach the personalized recommendation on member and non–member user. To address the member/non–member mode gap, Zhu et al.(Zhu et al., 2023) propose Model Extraction based MIA (ME-MIA) for sequential recommenders, which first extracts a surrogate model via black–box queries that align its recommended item set and rank order with the target using generic list–consistency objectives, and then performs membership inference using the surrogate’s richer signals (scores, ranks, similarities). ME-MIA targets sequential recommenders in a black-box setting by first extracting a surrogate model that imitates the target’s top – items and their rank order using two generic objectives—ranking consistency and positive-item consistency—derived solely from recommendation lists. Because the surrogate’s training does not reveal membership labels, a shadow model (trained with the same objective) is used to construct labeled data; rich signals from the surrogate (scores, ranks, similarities) then feed a binary classifier for membership inference. To reduce data demands, ME-MIA offers (i) a data-efficient variant that augments sequences by replacing actions with nearest-neighbor items in the surrogate’s embedding space, and (ii) a data-free variant that synthesizes sequences autoregressively and queries the target for soft labels, following data-free model-extraction practice. These steps yield effective and transferable attacks under black-box sequential settings. While ME-MIA demonstrates that membership inference is possible even in data–free settings, this comes with a loss in effectiveness; moreover, the role of item scoring in shaping attack success remains underexplored, and it is unclear whether ME-MIA readily generalizes to non-sequential RecSys. Despite recent progress at the user level, open challenges persist: how to design more efficient attacks, reduce confounding factors that obscure the signal (e.g., embedding dimensionality, top – , popularity effects), and accommodate the inherent heterogeneity of recommender systems with model–agnostic methods. Beyond traditional RecSys MIAs, He et al.(He et al., 2025c) (to our knowledge) present the first user–level attacks for LLM–based recommender systems, proposing five attacks—Inquiry, Hallucination, Semantic, Poisoning, and Contrast—that exploit LLMs’ memorization, text generation, and hallucination behaviors; evaluated across multiple popular LLMs, these attacks demonstrate clear effectiveness in revealing whether specific user were included in in–context prompts.
3.2. Interaction-level MIAs
@@
RecSys pose privacy risks beyond the user level: in addition to whether a user appears in the training set, one may care about the presence of a specific user–item interaction (interaction-level privacy). Yuan et al. (Yuan et al., 2023) conducted the first systematic study of interaction-level membership inference in federated recommender systems (FedRec), showing that such attacks are feasible and effective. In contrast to traditional FL-MIAs which have largely been designed for classification and aim to infer whether a sample was used for training (and sometimes which client used it), FedRecs are built on collaborative filtering, where item embedding dynamics can leak whether a client’s item set has changed. Moreover, each FedRec client corresponds to an individual user whose local data are non-IID, and the local dataset naturally consists of interacted and non-interacted items (only the former are meaningful targets for interaction-level MIAs), making the shadow-data assumptions commonly used in FL-MIAs impractical. Finally, unlike many FL classification settings that conceptually “share” a single global parameter set, FedRecs typically keep user embeddings private on the client while sharing (or aggregating) item embeddings on the server—an architectural choice that changes the attack surface and defenses. For these challenges, Yuan et al (Yuan et al., 2023) proposed a white-box interaction-level MIA on FedRec (c). IFed-MIA supposes the adversary know the target user’s updated public parameter (including gradient) and some hyperparameter like learning rate and negative sampling ratio. IFed-MIA observes only clients’ uploaded item-embedding updates and public hyperparameters (e.g., learning rate and negative-sampling ratio), but not private user embeddings or raw histories. From these uploads, the server can tell which items participated in a user’s local training; the problem then reduces to deciding whether a target user–item pair is a positive interaction. The attack relies on an empirical distance test: training three local models— on , on with a different private initialization, and on where is flipped—yields , enabling accurate inference of . Operationally, the server (i) identifies updated items, (ii) constructs candidate labels for them, (iii) trains a fake local model on the constructed dataset, and (iv) selects positives by comparing item-embedding distances between the uploaded parameters and the fake model, iterating until a prescribed ratio is met (Yuan et al., 2023). In practice, training hyperparameters such as the negative sampling ratio are typically not observable to an external adversary, casting doubt on the practicality of attacks that rely on them; even in federated recommenders, such hyperparameters are assumed known only to an honest-but-curious server that broadcasts them to clients. To improve interaction-level attacks beyond these assumptions, Zhong et al.(Zhong et al., 2024) propose a knowledge-graph (KG)–based method (MINER) that constructs a similarity matrix between recommended and interacted items using multiple distance measures and employs a cumulative-learning bilateral-branch network to emphasize long-tail items, thereby heightening risks tied to personalized interactions. However, while MINER targets the long-tail, it shows limited effectiveness in the low-FPR regime and its reliance on KGs constrains applicability to other RecSys architectures. Building on these observations, He et al.(He et al., 2025a) design a LiRA-based interaction-level attack (RecLiRA) that attains state-of-the-art true-positive rates at low false-positive rates across common RecSys models; moreover, they motivate a differential-privacy–inspired privacy score, , to quantify interaction-level risk and aggregate it to the user level(He et al., 2025b). Also, RecLiRA can be applied to all recommendation systems with implicit feedback. He et al.(He et al., 2025c) pioneer interaction–level MIAs for LLM–based recommender systems, exploiting LLMs’ tendency to memorize prompt content through direct inquiry and hallucination (“confusion”) attacks; experiments on several widely used LLMs and benchmark datasets show strong attack advantages, underscoring practical risks for in–context LLM RecSys.
3.3. Social-level MIAs
Unlike user-level and interaction-level membership inference attacks (MIAs), which only require knowing whether a user (or a user’s interactions) was included in the training set, Zhao et al.(Zhao et al., 2025) proposes a novel social-level privacy risk: inferring whether a link exists between two users. Given that traditional recommender systems generate recommendations based on a user’s own behavioural preferences and on the preferences of other users who are similar (homophily), extracting the socially relevant component from the mixed recommendation outputs is key to successfully recovering social ties. In this study we investigate whether an adversary—observing only recommendation results—can infer sensitive social relationships. The authors propose the SMIA framework, which incorporates an unsupervised attack based on recommendation similarity together with a supervised attack leveraging a shadow social graph and a dual-branch preference learning paradigm to disentangle social and behavioural preferences.
4. Evaluation
Building upon the aforementioned design principles, the evaluation of Membership Inference Attack (MIA) performance in recommender systems primarily centers on three fundamental metrics: the Area Under the Receiver Operating Characteristic Curve (AUC), the F1-score, and the True Positive Rate (TPR) to False Positive Rate (FPR) ratio. Each of these metrics serves a distinct yet complementary role in characterizing the attack’s discriminative capability, robustness, and real-world applicability. In this section, we present a comprehensive overview of the evaluation methodology, encompassing the benchmark datasets, representative recommendation models, and standardized performance indicators commonly adopted in empirical studies of MIAs on recommendation systems.
4.0.1. Datasets
MIAs methods on RecSys use the same datasets as other recommendation tasks. We list the widely used datasets and summarize the statistics in Table 1.
MovieLens The MovieLens dataset is one of the most widely adopted and benchmarked datasets in recommender system research. It comprises user–movie rating interactions and is available in multiple versions that differ in scale. The numeric suffix in each version denotes the approximate number of rating records it contains; for instance, MovieLens-1M includes about 1M user–item interactions.
Amazon The Amazon dataset comprises multiple domain-specific subsets categorized according to product types available on the Amazon platform. Each sub-dataset contains user reviews and metadata related to a particular product category. For instance, ADM, Beauty, Book, and Cell Phone represent the sub-datasets corresponding to Digital Music, Beauty products, books, and communication equipment, respectively.
LastFM The LastFM dataset is a music listening dataset collected from the Last.fm online music platform. It contains user–artist interaction records, such as listening histories and play counts, and is widely used for evaluating music recommendation and user preference modeling tasks.
Steam The Steam dataset was collected from Steam, one of the world’s largest digital distribution platforms for PC games. It contains detailed user–game interaction records, including transaction data such as game purchases and playtime durations. Among its various releases, Steam-200K is a widely adopted subset that serves as a benchmark for evaluating recommendation models in gaming-related domains.
Ta-feng The Ta-Feng dataset is a supermarket transaction dataset collected from a retail chain in Taiwan. It records detailed purchase histories, including user IDs, product IDs, quantities, and timestamps. Due to its sequential and temporal characteristics, it is widely used for sequential recommendation and next-item prediction research.
Yelp18 The Yelp dataset was originally compiled for the Yelp Dataset Challenge and contains users’ reviews of restaurants. The company Yelp3 is a platform that publishes crowd-sourced reviews of restaurants. which is a chance for students to conduct research or analysis on Yelp’s data and share their discoveries. A cleaned subset of Yelp reviews with user–business interactions and ratings, used for recommendation and review analysis.
Ciao The Ciao dataset was collected from the Ciao online product review platform, which allows users to rate, review, and socially interact with other users. It contains rich user–item interactions along with explicit social relations (e.g., trust networks), making it well-suited for evaluating social recommendation models. Each record includes user ratings, product metadata, and user trust links, enabling comprehensive analysis of both behavioral and social influence factors in recommendation tasks.
Flickr The Flickr dataset was extracted from the Flickr social image-sharing platform and includes users’ interactions with images (e.g., favorites, comments, or tags), as well as the underlying user–user social connections. It provides a representative benchmark for studying socially-aware recommendation systems, as it integrates both content-based and social network information, reflecting real-world scenarios where user preferences are shaped by social relationships and shared media interests.
| Dataset | User # | Item # | Interaction # |
| MovieLens-100K | |||
| MovieLens-1M | |||
| Yelp | |||
| Ta-Feng | |||
| Amazon-ADM | |||
| Amazon-Beauty | |||
| Amazon-Book | |||
| Amazon-Cell Phone | |||
| LastFM | |||
| Steam | |||
| Ciao | |||
| Flickr |
4.0.2. Models
To verify the effectiveness of the proposed methods, MIAs are often evaluated on various popular RecSys models. We list the widely used RecSys model structures as follows:
ICF calculates the similarity between items aiming to find the ones which are closed to users’ likes.
LFM builds a latent space to bridge user preferences and item attributes.
NCF A key collaborative filtering model that leverages neural network architectures.
LightGCN A state-of-the-art collaborative filtering model that simplifies graph convolution networks to enhance recommendation performance.
GRU4Rec is a session-based recommendation model that uses Gated Recurrent Units (GRUs) to model a user’s short-term click sequence and predict the next item. It’s typically trained with pairwise ranking losses (e.g., BPR/TOP1) and negative sampling, handling variable-length sessions efficiently.
BERT4Rec utilizes deep two-way transformer to model the sequence of user behavior, exhibiting excellent performance in multiple tasks.
STAMP not only captures the user’s general interests but also preserves the user’s current preference through a new short-term memory priority.
NARM consists of a global encoder and a local encoder. The latter encoder combined attention mechanism to attend large or small weights for different items.
CKE Collaborative Knowledge base Embedding integrates a knowledge graph into matrix-factorization by jointly learning user/item embeddings with KG structural/semantic regularizers (e.g., TransE/semantic embeddings), improving cold-start and representation quality.
KGAT Knowledge Graph Attention Network propagates user–item preferences over a KG with high-order neighbor aggregation via attention, then performs end-to-end recommendation with the KG-enhanced embeddings.
ECFAT Commonly cited as Explainable Collaborative Filtering with Attentive Transfer, it transfers item attribute/auxiliary information into CF via attention mechanisms to provide interpretable recommendations and better generalization
DiffNet++ A social recommendation model that iteratively diffuses user and item signals over the user–item bipartite and social graphs, capturing high-order social influence and preference propagation for improved link prediction.
DESIGN A social recommender that performs denoising/self-supervised learning on social graphs (and possibly user–item graphs) to reduce social noise and enhance representations before prediction.
GDMSR Graph Denoising based Multi-relational Social Recommendation leverages denoising objectives on social and interaction graphs to mitigate noisy/irrelevant relations while learning multi-relational user/item representations for recommendation.
4.0.3. Metrics